OpenStudy (anonymous):

Which function has a vertex at the origin? f(x) = (x + 4)2 f(x) = x(x – 4) f(x) = (x – 4)(x + 4) f(x) = –x2 @Michele_Laino

OpenStudy (michele_laino):

hint: the origin is the point \((0,0)\)

OpenStudy (anonymous):

i don't have no clue what this one is

OpenStudy (anonymous):

so it would be the last one right?

OpenStudy (michele_laino):

for example, if I replace \(x=0\) into the first equation, I get: \(f(0)=(0+4)^2=4^2=16\) so the point \((0,0)\) doesn't belong to that graph Yes! the last option

OpenStudy (anonymous):

like Michele_Laino said the orgin is 0,0 like on a graph